OpenStudy (anonymous):

Andre received $60 for his birthday and put it in his piggy bank. Each week, he puts 3 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 3w + 60, where w is the number of weeks he has saved 3 dollars. His sister Lana makes a table of reasonable values for w. Lana selects 12 as a value for w. What is the corresponding table entry for D? A. 372 B. 96 C. 75 D. 64

OpenStudy (anonymous):

I SAY 96, WHICH IS B DOES ANYONE ELSE AGREE? I WILL GIVE MEDAL AND FAN!

OpenStudy (anonymous):

d = 3w + 60 d = 3*12 + 60 d= 36 + 60 d = 96 correct?

OpenStudy (anonymous):

@MariahDimond @pooja195 @sissyedgar @dan815 @EclipsedStar

OpenStudy (anonymous):

i just need checking

OpenStudy (anonymous):

write the equation for a sec

OpenStudy (anonymous):

i just did

OpenStudy (anonymous):

up above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d = 3w + 60 d = 3*12 + 60 d= 36 + 60 d = 96 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mk